# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v3

@@ Abstract
This paper presents a novel methodology for determining representational structure, which builds upon the existing Spotlight Resonance method. This new tool is used to gain insight into how discrete representations can emerge and organise in autoencoder models, through a controlled ablation study in which only the activation function is altered. Using this technique, the validity of whether function-driven symmetries can act as implicit inductive biases on representations is determined. Representations are found to tend to discretise when the activation functions are defined through a discrete algebraic permutation-equivariant symmetry. In contrast, they remain continuous under a continuous algebraic orthogonal-equivariant definition. This confirms the hypothesis that the symmetries of network primitives can carry unintended inductive biases, which produce task-independent artefactual structures in representations. The discrete symmetry of contemporary forms is shown to be a strong predictor for the production of discrete representations emerging from otherwise continuous distributions — a quantisation effect. This motivates further reassessment of functional forms in common usage due to such unintended consequences. Moreover, this supports a general causal model for one mode in which discrete representations may form, and could constitute a prerequisite for downstream interpretability phenomena, including grandmother neurons, discrete coding schemes, general linear features and possibly Superposition. Hence, this tool and proposed mechanism for the influence of functional form on representations may provide insights into interpretability research. Finally, preliminary results indicate that quantisation of representations appears to correlate with a measurable increase in reconstruction error, reinforcing previous conjectures that this collapse can be detrimental.
@@ 1 Introduction
This introduction will begin by outlining how many of deep learning’s current functions are expected to influence the representations. This is hypothesised to constitute a function-driven emergent bias, motivating a comparative analysis against analogous functions adhering to differing symmetries. Finally, an approach is discussed for isolating such influences through a controlled ablation study.
@@ 1.1 Overview for How Functional Form Choices May Result in Biases
Directions singled out by a functional form’s structure may be termed a ‘privileged basis’ Elhage et al. (2022) or a set of ‘distinguished directions’ Bird (2025a). Such distinguished directions are commonplace through activation functions, normalisers, initialisers, regularisers, optimisers, architectures, operations, and more Bird (2025a). Hence, there are several avenues hypothesised through which representational influence may occur: directly through functional forms, indirectly through optimiser forms, and inherently through neuron connectivities. This proposed triad of influences would suggest that the representational alignment phenomenon may be systematic to deep learning’s anisotropic primitives and general practice. As a starting point, this work will isolate and determine the direct influence of activation functions as one of these causal influences on representations.
In particular, contemporary functional forms tend to make distinct the standard basis’ vectors, $\hat{e}_{i}$, and in such a way that each basis vector’s distinction is equal. This particular form of anisotropy can also be denoted through a permutation symmetry Godfrey et al. (2023). This can be promoted to a defining characteristic relation underlying contemporary functional forms and categorised accordingly Bird (2025a). This permutation symmetry is often due to their elementwise application, but can appear more generally. The specific relation depends on the primitive in question; for most activation functions, this is an algebraic permutation equivariance. This permutation equivariance is displayed in Eqn. 1111This is assuming a map $\rho:\mathcal{G}\rightarrow\mathrm{GL}_{n}\left(\mathbb{R}\right)$ representation of the group, in this case one which permutes the standard basis. This form of representation will be assumed and implicit in all groups represented as matrices moving forward.. The hypothesis is that the maximal symmetries to which functional forms belong are a strong predictor of emergent representational structure.
@@
The Spotlight Resonance Method Bird (2025b) has already empirically demonstrated how such activation functions can produce a tendency for representational alignments about the distinguished directions. This was achieved through transformations to the standard basis, and also altering the number of distinguished directions through the functional form. It was found that corresponding changes in the representations occurred, which followed the transformed structure of the distinguished directions. The conclusion was that functional forms stimulate this tendency towards representational alignment, producing emergent task-agnostic alignment structure in representations, and hence establishing a causal relationship between the two. The observed structure showed a tendency for a greater density of representations to cluster around the functional form’s distinguished directions. This work also indicated the presence of ’grandmother neurons’, demonstrating how many clusters represent a distinct semantic. However, it did not establish whether discretisation itself was a phenomenon of this choice of functional form.
@@ 1.2 Motivating The Approach for this Study
Isotropic Deep Learning Bird (2025a) takes the consequences of such observations a step further. The Spotlight Resonance Method determined that task-agnostic structure in representations is associated with a functional form’s distinguished directions. The Isotropic Deep Learning approach proposes that since this structure transforms with the functional form, the anisotropic functional form choices may be the fundamental causal mode that induces the array of interpretable structures to begin with.
In effect, anisotropic primitives provide an absolute frame, for each representation space, about which the representations organise and discretise into distinct clusters. Hence, it raises the question of whether the ubiquity of anisotropy in contemporary deep learning constitutes an underlying origin for many such interpretable and emergent-structure phenomena that are commonly observed. If this is the case, then it must be assessed whether this is beneficial or not. Furthermore, determining the validity of such a hypothesis would indicate whether these phenomena are truly fundamental to deep learning or stimulated by, and contingent on, the specific choices in functional form currently in use. This could have significant implications for the approach of network interpretability. Confirmation of this hypothesis would also demonstrate that such choices do carry inherent and unappreciated biases with their implementation.
This position motivates the comparative study of this work. Determining whether one can control and demonstrate the emergence and elimination of structure through broadening deep learning’s primitives to other forms, particularly those defined through differing maximal symmetries. The hypothesis is that each may provide its own inductive bias. One such proposed choice is by redefining primitives through an orthogonal algebraic-equivariance Bird (2025a), displayed in Eqn. 2. This is termed an ‘isotropic’ definition for activation function forms, and is a supergroup-symmetry of the original permutation symmetry, $S_{n}\subset\mathrm{O}\left(n\right)$.
@@
Choosing this definition for primitives is interesting, since it can eliminate all form-distinguished directions in a network. Therefore, functions developed to adhere to this modified symmetry definition can be used as a comparative mode to determine whether functional forms can induce these interpretable structures and indicate whether downstream alignment phenomena are attributable to anisotropic choices. If this is the case, it would demonstrate that these differing forms do carry unintended inductive biases, which would warrant major reevaluation of the forms used in standard practice.
1.3 Methodological Approach
@@
The Spotlight Resonance Method (SRM) Bird (2025b) hypothesised a hierarchy of anisotropic influences, particularly through the triad discussed earlier. These interacting distinguished bases may be highly non-trivial to predict the representational alignment of. Therefore, the causal mechanism was established in a clean, isolated and minimalistic way through ablation studies of only activation functions to minimise conflating or obscuring influences. Establishing this base case of a causal mechanism, from primitives to representations, can then be expected to carry downstream through more general networks built on these primitive influences. The SRM author’s approach of establishing such a causal mechanism in a minimalistic setting, which can then be expected to apply more generally, will be repeated in this methodology.
@@
Instead of rotating the privileged basis or altering the number of distinguished directions to determine alignment of representations, this study will compare representations between the presence and absence of any distinguished directions to determine if this is a causal mechanism for the production of approximately discrete representations.
@@
Due to the expected systematic nature of this inductive bias, care was taken to prevent the unintentional reintroduction of anisotropy. To achieve this, distinguished directions will be removed network-wide to avoid these confounding and obscuring interactions. This is achieved by utilising isotropic primitives and an autoencoder model for reconstruction  [^2] The use of an autoencoder model prevents anisotropy through choices in the output layer representation. For example, one-hot encoding in classification outputs is also a choice in form defined through permutation symmetry; therefore, it could induce further spurious structure. A reconstruction task using autoencoders may therefore mitigate this and enable more data-representative and ‘natural’ embeddings. Moreover, it has been speculated Bird (2025b, a) whether the neural-collapse classification phenomenon Papyan et al. (2020) is related to the proposed mechanisms. So autoencoder reconstruction provided a suitable minimalistic setting, free from these influences, to study form-induced representational quantisation., except for an activation function which can selectively reintroduce anisotropy through changes to its form. Hence, this isolates the causality of the phenomenon to the activation functions.
@@
This requires two activation functions, which are highly similar in action but differ in the presence or absence of anisotropy. The isotropic paper Bird (2025a) provides these preliminary isotropic activation functions, which enable this comparative analysis. An analogue of standard elementwise (anisotropic) tanh, displayed in Eqn. 3, can be compared with its orthogonal analogue Isotropic-Tanh, displayed in Eqn. 4 — each is displayed in multivariate form for comparison $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. These functions produce equivalent non-linear maps along the standard basis directions: $f_{\text{iso}}\left(\alpha\hat{e}_{i}\right)=f_{\text{aniso}}\left(\alpha\hat{e}_{i}\right)$ forall $\hat{e}_{i}$ and $\alpha$. Standard-Tanh is algebraically equivariant to the hyperoctahedral group (standard-basis permutation with sign-flips), $B_{n}$, whilst isotropic-tanh is algebraically equivariant to the orthogonal group, $\mathrm{O}\left(n\right)$, with $B_{n}\subset\mathrm{O}\left(n\right)$.
@@
$\mathbf{f}_{\text{aniso}}\left(\vec{x};\left\{\hat{e}_{i}\right\}_{\forall i}\right)=\sum_{i=1}^{N}\tanh\left(\vec{x}\cdot\hat{e}_{i}\right)\hat{e}_{i}$ (3)
@@
This enables a comparative ablation study of the influence of functional form symmetries on representations, and can determine a causal mechanism for the emergence of a discrete-like structure.
@@
This approach is also later extended to a Leaky-ReLU Maas et al. (2013) family, which are modified to adhere to the following symmetries $S_{n}\subset B_{n}\subset\mathrm{O}\left(n\right)$. This is discussed further in App. B.4 and indicates that algebraic symmetries are the primary predictor, since other analytical properties of Leaky-ReLU substantially differ from the tanh-like family of Eqns. 3 and 4.
This work also modifies the prior Spotlight Resonance Method (SRM). This modified tool is termed the Privileged-Plane Projective Method (PPP method), which operates on a similar principle. However, unlike SRM, this adaptation does not project out the vector length in the process. Hence, information regarding both angular and magnitude distributions is retained. This provides a more nuanced understanding of the network’s representational structures; however, SRM may still be preferable in determining statistical measures of the angular representations. Furthermore, this uniquely positions the Privileged-Plane Projective method as a tool for assessing whether neural refraction Bird (2025a) occurs for out-of-training-distribution outlier representations.
Additionally, the PPP method does not produce any angular-falloff smearing effect related to the cone width of SRM. Therefore, this modified method enables higher-fidelity angular maps of representations, which may be beneficial in many circumstances. The following section provides a detailed overview of this tool.
Using a comparative analysis between standard-Tanh and isotropic-Tanh, and similarly on Leaky-ReLU family of functions, this method can then determine whether there is a tendency for discrete representation clusters to form due to the task-agnostic functional form symmetries.
It indicates that algebraic permutation equivariance tends to produce a quantising inductive bias on otherwise continuous representations, rather than task-necessitated representational collapses. As a consequence, it suggests that downstream phenomena, such as interpretable neurons, are not fundamental but rather a direct consequence of anisotropic functional forms. This may have benefits in AI safety and interpretable networks; however, preliminary evidence also suggests that this task-agnostic collapse can be pathological, such as through reduced representational capacity.
1.4 Hypotheses of Function-Driven Discretisation
The primary hypothesis is that since isotropic functional forms are devoid of distinguished directions, they may consequently eliminate any tendency towards task-agnostic representational collapse aligned along these directions. This formed a prediction regarding why isotropic primitives may be beneficial in certain scenarios Bird (2025a), which is discussed further below. This motivates the importance of the present study as a starting point for probing this hypothesised representational collapse induced by primitives and establishing resultant consequences.
In particular, anisotropic forms may lead to a detectable tendency for representations to overly cluster around distinguished directions. This is due to the anisotropic functional form’s nature, where some angular arrangements may become overrepresented through optimisation due to their irregularity. This task-agnostic over-density bias may be detrimental, as representations may tend to collapse towards single directions, such as aligned or anti-aligned, and consequently lose informative representational degrees of freedom. This hypothesised loss in representational degrees of freedom may be damaging a network’s internal expressiveness. This quantising effect of anisotropy on representations is hypothesised also to affect the semantics being represented, resulting in a limited representational capacity and neural coding scheme. This may be exacerbated when the inputs and outputs are better represented continuously, and transformations between them should preserve an approximately continuous structure. Consequently, isotropic primitives are expected to remove this form-induced task-agnostic discretising bias on representations.
This may occur through several modes, hypothesised to stem from the underlying symmetry. As stated, particular symmetry-based constructions may produce distinguished directions that have more favourable maps for computations and representations shift accordingly.
Additionally, for general networks, their construction is not designed to vary predictably in response to group actions on representations. Therefore, if a network doesn’t commute with actions or remain invariant to them, then a shift in representations corresponding to some group action will then impact and interact non-trivially with the rest of the network. This is argued to produce a knock-on chain of interactions within the network. This influence is expected to result in unintended biases. It also highlights why groups/symmetry, rather than just a general action, may contribute an important characterising property, since many analogous actions within a group may each contribute a similar effect in the sub/space they transform and the resultant knock-on effects. Similarly, these biases may be expected to be largely consistent across sets of primitives that share the same maximal symmetry, due to the group structure of actions relating differing computational maps.
This is expected to differ from symmetries in end-to-end networks Bronstein et al. (2021), where their representations are designed to transform with the symmetry, resulting in the entire model often being either equivariant or invariant to its effect. Hence, those representative degrees of freedom connected through group actions become constrained by the group structure and, as a result, represent the pose information derived from the group. This is not the case in general networks, where a symmetry action does not pass cleanly through or identically cancel with layers. Therefore, these group-connected degrees of freedom may represent more general semantics and not indicate pose.
More informally, one can imagine that such symmetry actions can result in interactions within the network due to its structure and consequently alter dynamics, rather than passing straight through unimpeded by design. In this informal heuristic, a layer may ‘absorb’ a symmetry action in a non-invariant/equivariant manner, or have an interaction with its differing symmetry. This may mean the symmetry does, in principle, ‘act’ on the network and hence produce a realisable and tangible effect. This suggests that the maximal symmetry may be an important consideration in general networks, resulting in similar inductive biases, due to function-driven biases. These are felt to be underappreciated in significance.
One can then consider permutation-like symmetries about a particular basis. These result in an orthant partitioning of the space. Depending on the particular permutation symmetry, sets of these orthants may be analytically degenerate, producing transformed copies of equivalent maps — a manifestation of the underlying group definition. The network’s optimisation may then shape representations about these orthants of differing and equal mapping, to achieve the desired computation. This may result in representations that organise differently over these maps related by the symmetry of the functional form. In their extreme, they may cluster densely over particular preferred regions, producing a quantisation effect on representations. This is considered a general mode through which function-driven symmetries may influence representations in permutation-defined forms — a response of optimisation to the symmetry. One can then compare this to an analogous function, contingent upon a non-permutation defined form, allowing for a comparison of their respective biases. This general approach will be used to study how various function-driven symmetries interact and act on the network and its representations, determining the various biases induced.
This mechanism would aid in explaining why the SRM study Bird (2025b) detected axis-aligned structure. Incidental alignment with directions chosen to be defined and observed as ‘individual neurons’ would otherwise have a vanishingly small likelihood of occurring in high-dimensional spaces. This is because there are vastly more orientations that do not align with the standard basis than those that do in high-dimensional spaces. However, if the analytic properties of permutation-based functions produce maps which differ along these directions, ‘distinguished directions’, then the symmetry breaking of the space would single out such directions and make alignment less improbable. Therefore, it indicates an unintended bias of the model. A comparison between anisotropic and isotropic forms can directly probe such questions by adding and removing any distinction between directions.
Quantisation is argued to have potential pathological effects on the network. One heuristic is that the underlying distributions might be better represented continuously, as this aligns with the frequently continuous real-world semantics being represented, such as broad object morphologies and arrangements, multivariate signals, lighting and shades, continuous expressions, and modelling distributions, etc. Therefore, enabling latent variables and representations to be continuous and potentially more interpolatable at various levels through a network, instead of an imposed task-agnostic structure, may be beneficial for many tasks. This may increase representational capacity and organisation of the representations.
This connects to the coding scheme of the network. Axis-aligned representations, previously observed to frequently emerge Bau et al. (2017); Elhage et al. (2022); Bird (2025b), approximate a local coding scheme Sherrington (1940); Konorski (1968); Barlow (1953); Gross (2002); Connor (2005) for the network. This is where one neuron corresponds to one semantic. This coding scheme has decreased concept-interference, but has a much lower representational capacity than other coding schemes. Models have been demonstrated to produce other interference-capacity balances with various input sparsities Elhage et al. (2022); Bird (2025b), suggesting that axis-alignment is not an appropriate universal inductive bias — further motivating the study of alternative functional form inductive biases and their effects on representations.
A functional form that does not distinguish particular directions may therefore not deter such continuous and interpolatable structures present in the dataset, and this may be advantageous to some applications. Yet, sometimes discretisation may be appropriate, such as in classification networks, where abstracting away redundant detail can be beneficial — though in autoencoding models this would not be an expectation. In any case, rigidly imposing this through a task-agnostic, function-driven, inductive bias about a specific arrangement might be detrimental. For example, when reducing the dimensionality of a dataset inherently containing discrete representations via a linear layer, to some arbitrary lower dimension, one would not expect it to be always appropriate for those discrete directions to be best represented through an alignment to an orthonormal basis or other imposed geometry for that space — as these are as arbitrary to the bottleneck chosen.

### v4

@@ Abstract
Presented is a novel methodology for determining representational structure, which builds upon the existing Spotlight Resonance method. This new tool is used to gain insight into how discrete representations can emerge and organise in autoencoder models, through a controlled ablation study that alters only the activation function. Using this technique, the validity of whether function-driven symmetries can act as implicit inductive biases on representations is determined. Representations are found to tend to discretise when the activation functions are defined through a discrete algebraic permutation-equivariant symmetry. In contrast, they remain continuous under a continuous algebraic orthogonal-equivariant definition. This confirms the hypothesis that the symmetries of network primitives can carry unintended inductive biases, leading to task-independent artefactual structures in representations. The discrete symmetry of contemporary forms is shown to be a strong predictor for the production of symmetry-organised discrete representations emerging from otherwise continuous distributions — a quantisation effect. This motivates further reassessment of functional forms in common usage due to such unintended consequences. Moreover, this supports a general causal model for a mode in which discrete representations may form, and could constitute a prerequisite for downstream interpretability phenomena, including grandmother neurons, discrete coding schemes, general linear features and a type of Superposition. Hence, this tool and proposed mechanism for the influence of functional form on representations may provide insights into interpretability research. Finally, preliminary results indicate that quantisation of representations correlates with a measurable increase in reconstruction error, reinforcing previous conjectures that this collapse can be detrimental.
@@ 1 Introduction
This introduction will begin by outlining how many of deep learning’s current functions are expected to influence the representations. This is hypothesised to constitute a function-driven emergent bias, motivating a comparative analysis against analogous functions that adhere to differing symmetries. Finally, an approach to isolate such influences through a controlled ablation study is discussed. This is centred around empirically testing a prediction made for such functional forms [4].
@@ 1.1 Overview for How Functional Form Choices May Result in Biases
Directions singled out by a functional form’s structure may be termed a ‘privileged basis’ [8] or a set of ‘distinguished directions’. Such distinguished directions are commonplace through activation functions, normalisers, initialisers, regularisers, optimisers, architectures, operations, and more [4]. Hence, there are several avenues hypothesised through which representational influence may occur: directly through functional forms, indirectly through optimiser forms, and inherently through neuron connectivities. This proposed triad of influences suggests that the representational alignment phenomenon may be systematic to deep learning’s anisotropic primitives and general practice. As a starting point, this work will isolate and determine the direct influence of activation functions as one of these causal influences on representations.
In particular, contemporary functional forms tend to make distinct the standard basis’ vectors, $\hat{e}_{i}$, and in such a way that each basis vector’s distinction is equal. This particular form of anisotropy can also be denoted through a permutation symmetry [10], represented standardly. This can be promoted to a defining characteristic relation underlying contemporary functional forms and categorised accordingly [4]. This permutation symmetry is often due to their elementwise application, but can appear more generally. In effect, anisotropic primitives provide an ‘absolute frame’, for each representation space, about which the representations may organise and begin discretising into more distinct symmetry-led clusters.
The specific relation depends on the primitive in question; for most activation functions, this is an algebraic permutation equivariance, displayed in Eqn. 1111This is assuming a map $\rho:\mathcal{G}\rightarrow\mathrm{GL}_{n}\left(\mathbb{R}\right)$ representation of the group, in this case one which permutes the standard basis. This form of representation will be assumed and implicit in all groups represented as matrices moving forward.. Where $S_{n}$ is the permutation group, $\mathbf{P}$ is an element of that group represented standardly as a matrix, and $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the function of interest. The hypothesis is that the maximal symmetries to which functional forms belong are a strong predictor of the emergent, representational structure organised by symmetry.
@@
The Spotlight Resonance Method [5] has already empirically demonstrated that such activation functions can produce a tendency toward representational alignments about the distinguished directions that transform with this basis, thereby inferring an absolute frame. However, this study did not demonstrate that these functions were sufficient to produce these phenomena.
This was achieved through transformations to the standard basis and also varying the number of distinguished directions through the functional form. The former amounts to different matrix representations of the permutation group influencing representations, and the latter often corresponds to differing discrete rotation groups which vary in rotation angle (although some notably do not form group-defined forms).
It was found that corresponding changes in the representations occurred, which followed the transformed structure of the distinguished directions. The conclusion was that these functional forms stimulate the tendency towards representational alignment, producing emergent task-agnostic alignment structure in representations.
This established a causal relationship between the two; both transformed in step, and differing discrete symmetries produced similarly organised clusters. Hence, the observed structure showed a tendency for representations to cluster more densely around the functional form’s distinguished directions. This work also indicated the presence of ’grandmother neurons’, demonstrating how many clusters represent a distinct semantic.
However, this SRM study [5] did not establish whether discretisation itself was a phenomenon of this choice of symmetry-defined functional form, as it did not compare continuous symmetry-defined forms against discrete symmetry-defined forms. This sufficiency is determined in the present work.
@@ 1.2 Motivating The Approach for this Study
Isotropic Deep Learning [4] takes the consequences of such observations a step further. The Spotlight Resonance Method determined that task-agnostic structure in representations is associated with a functional form’s distinguished directions. The Isotropic Deep Learning approach proposes that, since this structure observably transforms with the group’s representation, which defines the functional form, then the choice of an anisotropic form may be the fundamental causal mode that induces the array of interpretable structures to begin with.
This work directly explores this prediction: testing whether downstream alignment is predicated on anisotropic forms. If validated, it reveals a systematic, unappreciated foundational bias embedded in modern deep learning. Additionally, it enables a mode of control over this bias and the resulting generation of function-induced structure. If this prediction is validated, it is singularly pivotal to assess whether this unintended bias is beneficial or detrimental — preliminary results of App. D support the latter. It would also have significant ramifications for the interpretability of networks, where the appearance and organisation of discrete-like features can be controlled.
Hence, it raises the question of whether the ubiquity of anisotropy in contemporary deep learning constitutes the underlying origin of many such interpretable and emergent structural phenomena commonly observed. If these are symmetry-led, then the organisation will be determined by the symmetry.
This position motivates the comparative study of this work. Determining whether one can control and demonstrate the emergence and elimination of organised structure through broadening deep learning’s primitives to other forms, particularly those defined through differing maximal symmetries.
The hypothesis is that each may provide its own inductive bias. One such proposed choice is by redefining primitives through an orthogonal algebraic-equivariance [4], displayed in Eqn. 2 for the orthogonal group $\mathrm{O}\left(n\right)$ and its elements represented standardly as matrices $\mathbf{R}$. This is termed an ‘isotropic’ definition for activation function forms, and is a supergroup-symmetry of the original permutation symmetry, $S_{n}\subset\mathrm{O}\left(n\right)$.
@@
Choosing this definition for primitives is interesting, since it can eliminate all form-distinguished directions in a network. Therefore, functions developed to adhere to this modified symmetry definition can serve as a comparative mode to determine whether functional forms can induce these interpretable structures and to indicate whether downstream alignment phenomena are attributable to anisotropic choices.
If this is the case, it would demonstrate that these differing forms do carry unintended inductive biases, which would warrant major reevaluation of the forms used in standard practice.
The investigation of this mechanism improves upon the SRM [5] by explaining why axis-aligned structures are produced, not just how they behave under basis changes.
Incidental alignment with directions chosen to be defined and observed as ‘individual neurons’ would otherwise have a vanishingly small likelihood of occurring in high-dimensional spaces. This is because there are vastly more orientations that do not align with the standard basis than those that do in high-dimensional spaces. However, if the analytic properties of permutation-based functions produce maps which differ along these directions, ‘distinguished directions’, then the symmetry breaking of the space would single out such directions and make alignment less improbable.
Therefore, it indicates an unintended bias of the model. A comparison between anisotropic and isotropic forms can directly probe such questions by adding or removing any distinction between directions.
Consequently, this study can be considered to determine a more foundational prerequisite phenomenon than the Spotlight Resonance Method observed and causally explained. This prior method elucidated how functional forms can causally influence alignment structure. In contrast, this work investigates whether forms can lead to the emergence of any induced task-agnostic structure. Determining whether current forms induce representational quantisation may indicate a causal mechanism for these other downstream phenomena. This then indicates whether these phenomena are truly fundamental to deep learning or are stimulated by, and hence contingent on, the specific choices in functional form currently in use. This could have significant implications, demonstrating that such choices do carry inherent, unappreciated biases in their implementation and affect deductions made in the approach to network interpretability.
1.3 Discussion of Methodological Approach
@@
The Spotlight Resonance Method (SRM) hypothesised a hierarchy of anisotropic influences, particularly through the triad discussed previously. These interacting distinguished bases may be highly non-trivial to predict any symmetry-led organisation of representations. Therefore, the causal mechanism was established in a clean, isolated, and minimalistic way through ablation studies of only the activation functions, to minimise confounding or obscuring influences.
Establishing this foundational case of a causal mechanism, linking symmetry-defined primitive functional forms to a form of representation influence, may then be expected to carry implications downstream through more general networks built on these primitive influences, where its more complicated interactions may be further predicted and studied. A discussion of generalising the predictions to other architectures is discussed in App. C.1. Hence, the SRM approach of establishing such a causal mechanism in a minimalistic setting will be repeated in this methodology.
@@
As stated, instead of rotating the privileged basis or altering the number of distinguished directions to determine the alignment of representations, this study will compare representations in the presence and absence of any distinguished directions to determine whether this is a causal mechanism for the production of approximately discrete representations.
@@
Due to the expected systematic nature of this inductive bias, great care is necessitated to prevent the unintentional reintroduction of anisotropy. To achieve this, distinguished directions will be removed network-wide to avoid these confounding and obscuring interactions. This is achieved by utilising isotropic primitives and an autoencoder model for reconstruction [^2] Additionally, the use of an autoencoder model is essential to prevent anisotropy reintroduced through choices in the output layer representation. For example, one-hot encoding in classification outputs is also a choice in form defined through permutation symmetry; therefore, it could induce further spurious structure. A reconstruction task using autoencoders may therefore mitigate this and enable more data-representative and ‘natural’ embeddings. Moreover, it has been speculated [5, 4] whether the neural-collapse classification phenomenon [21] is related to the proposed mechanisms through this permutation-defined classification layer. So, autoencoder reconstruction provided a suitable minimalistic setting, free from these influences, to study form-induced representational quantisation directly and without confounders., except for an activation function which can selectively reintroduce anisotropy through changes to its form. Hence, this isolates the phenomenon’s causality to the activation functions.
This requires two activation functions that are highly similar in action but differ only in whether anisotropy is present.
The Isotropic paper [4] provides these preliminary isotropic activation functions, which enable this comparative analysis.
@@
An analogue of standard elementwise (anisotropic) tanh, displayed in Eqn. 3, can be compared with its orthogonal analogue Isotropic-Tanh, displayed in Eqn. 4 — each is displayed in multivariate form for comparison $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Standard-Tanh is algebraically equivariant to the hyperoctahedral group (standard-basis permutation with sign-flips), $B_{n}$, whilst isotropic-tanh is algebraically equivariant to the orthogonal group, $\mathrm{O}\left(n\right)$, with $B_{n}\subset\mathrm{O}\left(n\right)$. The term $\left\|\vec{x}\right\|=\left\|\vec{x}\right\|_{2}$ is the standard two-norm of the vector $\vec{x}\in\mathbb{R}^{n}$.
@@
$\mathbf{f}_{\text{aniso}}\left(\vec{x};\left\{\hat{e}_{i}|\forall i\right\}\right)=\sum_{i=1}^{N}\tanh\left(\vec{x}\cdot\hat{e}_{i}\right)\hat{e}_{i}$ (3)
@@
These functions produce exactly equivalent non-linear maps along the standard basis directions: $f_{\text{iso}}\left(\alpha\hat{e}_{i}\right)=f_{\text{aniso}}\left(\alpha\hat{e}_{i}\right)$ forall $\hat{e}_{i}$ and $\alpha$. They only diverge away from the standard basis, but remain highly analogous in their non-linear map.
@@
This enables a comparative ablation study of the influence of functional-form symmetries on representations and is sufficient to determine a causal mechanism for the emergence of a discrete-like structure.
@@
Similar to this implementation, experiments are later extended to a Leaky-ReLU [18] family, which are modified to adhere to the following symmetries $S_{n}\subset B_{n}\subset\mathrm{O}\left(n\right)$. This is discussed further in App. B.4 and indicates that algebraic symmetries are the primary predictor, since other analytical properties of Leaky-ReLU substantially differ from the tanh-like family of Eqns. 3 and 4.